Suppose you flip a coin 3 times. What is the probability that exactly 2 of your flips were heads?
Answer by stanbon(75887)

Suppose you flip a coin 3 times. What is the probability that exactly 2 of your flips were heads?
----
Binomial Problem with n = 3 ; p(head) = 1/2
---
P(x = 2) = 3c2(1/2)^2*(1/2) = 3/8
===================================
Cheers,
Stan H.
===========